BOUND ROTATION–VIBRATION STATES FOR THE Ne_{2}XH (Ã^{2}Σ, v_{HX} = 0, X=O,S) VAN DER WAALS COMPLEXES

\underline{HEE-SEUNG LEE}, AND ANNE B. MCCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The rotation–vibration states of Ne_{2}XH (Ã^{2}Σ) are obtained by variational calculations. The potential energy surfaces are described by a sum of the Ne–HX empirical potential surface developed by us^{a}^{b} and the Ne–Ne potential of Aziz.^{c} In these calculations, the orthogonalized satellite coordinates are used in order to properly describe both the T–shaped and linear geometries of the complexes. The ground states of the Ne_{2}XH complexes are localized in the region of the potential that corresponds to the T–shaped geometry with the hydrogen atom pointing toward the center of mass of neon dimer. In the case of Ne_{2}OH, there are low lying states that are localized in the region of the potential that corresponds to the linear Ne–OH–Ne geometry. The complete rotation-vibration energy level progressions of Ne_{2}OH and Ne_{2}SH are presented. Comparisions with available experimental data and previous diffusion Monte Carlo results are also made.

aC. C. Carter, T. A. Miller, H. -S. Lee, A. B. McCoy and E. F. Hayes J. Chem. Phys \underline{110}, 5065 (1999)

bH. -S. Lee, A. B. McCoy, L. B. Harding, C. C. Carter and T. A. Miller J. Chem. Phys \underline{111}, 10053 (1999)

cR. A. Aziz and M. J. Slaman Chem. Phys \underline{130}, 187 (1989)